OpenStudy (anonymous):

how do you solve (√x+5)+(√x-2)=5

OpenStudy (anonymous):

Rather, \[\sqrt{x}+\sqrt{x}=2\sqrt{x}\]

OpenStudy (anonymous):

@BangkokGarrett as for this question, how can we apply it? how do we solve it!??

OpenStudy (anonymous):

@BangkokGarrett i think it has something to do with quadratic equation... please help!

OpenStudy (anonymous):

I'll do the first couple of steps \[\sqrt{x}+5+\sqrt{x}-2=5\] \[2\sqrt{x}+3=5\] Just keep going 'til you have the sqrt(x) isolated on the left of the equation

OpenStudy (anonymous):

No need for the quadratic equation here.

OpenStudy (anonymous):

@BangkokGarrett but i have a question, then don't you have to multiply 5 and 2 on the other side as well, meaning: √+5+x−√−2=5x5x2 ?